OpenStudy (anonymous):

(x + 2y)(3x − 4y + 5)

OpenStudy (kc_kennylau):

(x+2y)(3x-4y+5) =x(3x-4y+5)+2y(3x-4y+5) =3x^2-4xy+5x+6xy-8y^2+10y =3x^2+2xy-8y^2+5x+10y

OpenStudy (anonymous):

Okay I still kind of don't understand that.

OpenStudy (anonymous):

Oh I see now. Thank you.

OpenStudy (kc_kennylau):